Sasha owns two investments, A and B, that have a combined total value of 41,600 dollars. Investment A is expected to pay 21,700 dollars in 1 year(s) from today and has an expected return of 12.79 percent per year. Investment B is expected to pay 42,573 dollars in T years from today and has an expected return of 6.96 percent per year. What is T, the number of years from today that investment B is expected to pay 42,573 dollars? Round your answer to 2 decimal places (for example, 2.89, 14.70, or 6.00).
Answer #1
Investment A value at Y1=                21,700.00
Expected return= 12.79%
Investment A value today= 21700/(1+12.79%)
Investment A value today=                19,239.29
Funds available today                41,600.00
Investment A              (19,239.29)
Investment B                22,360.71
Amount                42,573.00
Return 6.96%
time N
42573= 22360*(1+6.96%)^N
42573/22360= (1+6.96%)^N
                                          1.90 (1+6.96%)^N
log 1.90398 N * Log 1.0696
0.27966= N* 0.02922139
N= 0.27966/0.02922139
N=                          9.57